The number of disintegrations per second of C14 nuclei in a 1 milligram sample from a dead tree is 85.3% of that of the disintegrations of 1 milligram sample from a living tree. What year did the tree die?
Answer #1

Decay constant of C-14 = 3.84 * 10-12 s-1

We have the relation: A = A0 * e-kt

Where A is the current activity (0.853A0), A0 is the initial activity, k is the decaty constant and t is the time

Using the given values, 0.853 * A0 = A0 * e-3.84 * 10-12 * t

Solving: 0.159 = 3.84 * 10-12 * t

t = 4.141 * 1010 s = 1312.98 years, therefore the tree dies 1313 years ago menaing in the year 717 AD